1. (i) Prove that if m and n are integers and mn is even, then m is even or n is even.
(ii) Show that if n is an integer and n3 + 5 is odd, then n is even using
(a) a proof by contraposition
(b) a proof by contradiction
(iii) Prove that if n is an integer and 3n + 2 is even, then n is even using
(a) a proof by contraposition
(b) a proof by contradiction
(iv)ProvethepropositionP(0),whereP(n)istheproposition”ifnisapositiveintegergreater than 1, then n2 > n.” What kind of proof did you use?
(v) Prove the proposition P(1), where P(n) is the proposition ” If n is a positive integer, then n2 ≥ n.” What kind of proof did you use?
1
Expert's answer
2020-06-22T17:41:23-0400

1.(i) Suppose "mn" is even. Then we can choose an integer "k" such that "mn=2k." If "m" is even, then there is nothing more to proove, so suppose "m" is odd. Hence "m=2j+1" for some integer "j." We get


"(2j+1)n=2k""2jn+n=2k""n=2(k-jn)"

Since "k-jn" is an integer, it follows that "n" is even.


(ii)

a) a proof by contraposition

The contrapositive is “If "n" is odd, then "n^3+5" is even.” Assume that "n" is odd. We can now write "n=2k+1" for some integer "k." Then


"n^3+5=(2k+1)^3+5=8k^3+12k^2+6k+1+5=""=8k^3+12k^2+6k+6=2(4k^3+6k^2+3k+3)."

Thus "n^3+5" is two times some integer, so it is even by the definition of an even integer.

b) a proof by contradiction

Suppose that "n^3+5" is odd and that "n" is odd. Since "n" is odd, the product of odd numbers is odd. So we can see that "n^3" is odd. But if we subtract "5," then the difference between the two odd numbers "n^3+5" and "n^3" is even. Thus, our assumption was wrong and it is a contradiction. 


Therefore if "n" is an integer and "n^3+5" is odd, then "n" is even.


(iii)

(a) a proof by contraposition 

(Contrapositive: If "n" is even, then "3n+2" is even) Suppose that the conclusion is false, i.e., that "n" is even. Then "n=2k" for some integer "k."

Then "3n+2=3(2k)+2=6k+2=2(3k+1)."

Thus "3n+2" is even, because it equals "2j" for an integer "j=3k+1." So "3n+2" is not odd. 

b) a proof by contradiction

Assume that the conclusion is false, i.e., that "n" is even, and that "3n+2" is odd. Then "n=2k" for some integer "k" and "3n+2=3(2k)+2=6k+2=2(3k+1)."

Thus "3n+2" is even, because it equals "2j" for an integer "j=3k+1."This contradicts the assumption “"3n+2" is odd”. This completes the proof by contradiction, proving that if "3n+2" is odd, then "n" is odd. 


(iv)

Note that "P(0)" is that "If is a positive integer greater than "1", then "0^2>0" ".

Since " is a positive integer greater than "1" " is false the the overall statement is true.

This is a vacuous proof.


(v)

Note that "P(1)" is that "If "1" is a positive integer, then "1^2\geq1"".

Since "1^2=1," then "P(1)" is true.

This is a trivial proof.